This is proving a trigonemetric function.
Verify (1+sec(-x))/(sin(-x)+tan(-x)=-csc(x)
So far I have: (1-(1/cos(x)))/(-sin(x)-(sin(x)/cos(x))= -1/sin
I think the numerator might equal sin(x) but I'm not too sure and the denomintator confuses me. Please help.
Thanks!

Answer by Alan3354(69443)

This is proving a trigonemetric function.
Verify (1+sec(-x))/(sin(-x)+tan(-x)=-csc(x)
-csc(x) = csc(-x), so
All angles are -x
-------------------
1+sec = csc*(sin + tan) = 1 + (1/sin)(sin/cos)
1+sec = 1 + 1/cos
1+sec = 1+sec